### v2

# A prime number “game of life”: can $\lfloor y\cdot p\#\rfloor$ be prime for all $p\geq 2$?

Abstract.
A new sequence in the spirit of the Mills primes is presented and its properties are investigated.
1. Introduction
In the 1940s, W. H. Mills [8] examined prime-representing functions, most notably he pointed out, based on a result of Ingham [5], that there exists a constant $A$ such that $\lfloor A^{3^{n}}\rfloor$ is prime for all integers $n\geq 1$. Mills’ ansatz reproduces only a very sparse subset of primes. There are several variations to obtain a larger subset of primes, e. g. by Elsholtz [2], who conjectures the existence of a constant $A$ such that $\lfloor A^{(n+1)^{2}}\rfloor$ is prime for all $n\geq 1$; Kuipers [6], or Plouffe [12], who consider sequences of the form $\lfloor A^{c^{n}}\rfloor$ for $c$ as small as 1.01.
In this paper, we give an even more slowly growing formula—in fact, for its kind as an asymptotically exponentially growing function, it exhibits the slowest possible growth rate while still producing prime numbers, and a comparatively large subset at that.
While large portions of this paper consist of artisan work with well-known tools of number theory, there may something to be gained in-between the lines, for example a way to calculate Chebyshev’s $\theta(x)$ directly from known values of $Li(x)-\pi(x)$ which we haven’t seen before in literature. There are also a couple of open problems which may or may not be worth investigating further.
2. Definitions
$p,q$ variables for prime numbers
$p_{n}$ the $n^{th}$ prime; ${p_{n}}\in\{2,3,5,7,11,13,17\ldots\}$
$\pi(x)$ the prime counting function; the number of primes $p_{n}\leq x$
p#p\# the primorial function: $p\#=2\cdot 3\cdot 5\cdot 7\cdot 11\cdot\ldots\cdot p$
1# = 1 by convention
$\Phi(n,p)$ number of integers $\leq n$ that are relatively prime to p#p\#
$\log x$ natural logarithm of $x$
$\lfloor x\rfloor$ the floor function; the largest integer $\leq x$
$[a,b[$ the open interval containing real numbers $x$ such that $a\leq x<b$
$|x|$ absolute value of $x$; |x|=x|x|=-x for $x<0$
3. Setting the stage
[…]
Lemma: Any real number $y$ such that $q=\lfloor y\cdot p\#\rfloor$ is a prime number for every $p\geq 2$ is contained in the interval $[1,1.5[$.
If $y<1$, then $\lfloor y\cdot 2\#\rfloor<2$, which is less than the first prime number.
For $y\geq 1.5$, $\lfloor y\cdot 2\#\rfloor$ must be an odd prime 2m12m-1 for some integer $m>1$. Then $\lfloor y\cdot 3\#\rfloor$ is a number in the interval [6m3,6m1][6m-3,6m-1], only the latter of which can be prime because 6m36m-3 is divisible by 3 and 6m26m-2 is divisible by 2. Likewise, any integer in the successive intervals $p\#\cdot m-p,p\#\cdot m-1]$ is divisible by a small prime $\leq p$ except for $p\#\cdot m-1$.
Hence there would have to be an integer $m$ for which $p\#\cdot m-1$ is always prime, i.e. the exponentially growing sequence f(0)=m1f(0)=m-1, $f(s+1)=(f(s)+1)\cdot p_{s}-1$ must only produce primes for all $s\geq 1$, which is practically impossible (though we have not seen this being explicitly disproven yet!).
It is possible, however, to find arbitrarily long finite sequences for some $m$:
Table 1. Smallest integers $m$ for which $p\#\cdot m-1$ is prime for $p\leq r$. The first unknown entry for $r=47$ has $m>1.6\cdot 10^{12}$ and is expected to occur when $10^{14}<m<10^{15}$.
$r$ $m$ $r$ $m$
5 2 23 384427
7 2 29 16114470
11 9 31 259959472
13 9 37 13543584514
17 9 41 100318016379
19 224719 43 100318016379
To find out about an admissible value for $y$, we have to look for primes $q$, where each $q$ requires the primality of $\lfloor q/p\#\cdot r\#\rfloor$ for every $r\leq p$. In the following setup, $Y$ will denote the interval in which any number $y$ as described above can lie, setting $Y=[1,1.5[$ for a start, and then successively sharpen the bounds where, for every $p$, $Y\cdot p\#$ must contain at least one prime to satisfy the conditions.
$\bullet\,$ Stage 1: $p=2\,\rightarrow\,2\cdot Y=[2,3[\,\rightarrow\,\lfloor 2\cdot Y\rfloor=2$ – no further action required yet
$\bullet\,$ Stage 2: $p=3\,\rightarrow\,6\cdot Y=[6,9[$ – the only prime in here is 7, so Y narrows down to $\frac{7}{6},\frac{8}{6}[$
$\bullet\,$ Stage 3: $p=5\,\rightarrow\,30\cdot Y=[35,40[$ – one prime (37) in this interval, $Y$ shrinks to $\frac{37}{30},\frac{38}{30}[$
$\bullet\,$ Stage 4: $p=7\,\rightarrow\,210\cdot Y=[259,266[$ – once again, one prime in here and any possible $y$ lies between $\frac{263}{210}$ and $\frac{264}{210}$
$\bullet\,$ Stage 5: $p=11\,\rightarrow\,2310\cdot Y=[2893,2904[$ – now it’s getting interesting, since there are two primes in that interval, 2897 and 2903. We will split $Y$ into two separate intervals, $Y_{1}=[\frac{2897}{2310},\frac{2898}{2310}[$ and $Y_{2}=[\frac{2903}{2310},\frac{2904}{2310}[$ and proceed
$\bullet\,$ Stage 6: $p=13\,\rightarrow\,$ each $Y_{1}$ and $Y_{2}$ brings forth one prime, 37663 and 37747 respectively. Cutting back both $Y$’s, then
$\bullet\,$ Stage 7: $p=17\,\rightarrow\,$ $Y_{1}$ leads to 640279, whereas $Y_{2}$ gives 641701 and 641713, and $Y_{2}$ will be split into $Y_{2}$ and $Y_{3}$ in accordance with the primes
$\bullet\,$ Stage 8: $p=19\,\rightarrow\,9699690\cdot Y_{1}$ contains a prime triplet, 12165311+$d$ for $d=\{0,2,6\}$, while $Y_{2}$ and $Y_{3}$ both fail to produce any further prime
$\bullet\,$ etc.
The results can be put into a nice grid:
Table 2. This seems to be the new “game of life”, a family tree of prime numbers. Prime triplets $q+\{0,2,6\}$ or $q+\{0,4,6\}$ like in stage 8 are rather rare in the process, one finds the next example at stage 152 ($p$ = 881), and there are no prime quadruplets $q+\{0,2,6,8\}$ in any stage $\leq$ 331.
$\lfloor Y\rfloor$ 1
$\cdot 2$ +0
$\cdot 3$ +1
$\cdot 5$ +2
$\cdot 7$ +4
$\cdot 11$ +4 +10
$\cdot 13$ +2 +8
$\cdot 17$ +8 +2 +14
$\cdot 19$ +10 +12 +16
$\cdot 23$ +16 +8 +16 +18
$\cdot 29$ +16 +6 +14 +24 +26 +10
$\cdot 31$ +14 +24 +6 +14
$\cdot$... +… +… +…
Will the evolution go on forever? The following table shows how many partial intervals (or primes, respectively) $n$ there are after stage $s$, up to $s=60$:
Table 3. Talking about a bottleneck: the only prime at stage 29 is 350842542483891235293716663559065020274899073 ($\approx 3.5\cdot 10^{44}$).
$s$ $p$ $n$ $s$ $p$ $n$ $s$ $p$ $n$
1 2 1 21 73 5 41 179 18
2 3 1 22 79 4 42 181 17
3 5 1 23 83 6 43 191 14
4 7 1 24 89 3 44 193 24
5 11 2 25 97 2 45 197 24
6 13 2 26 101 1 46 199 28
7 17 3 27 103 3 47 211 30
8 19 3 28 107 1 48 223 36
9 23 4 29 109 1 49 227 49
10 29 6 30 113 3 50 229 44
11 31 4 31 127 2 51 233 52
12 37 5 32 131 5 52 239 53
13 41 5 33 137 6 53 241 55
14 43 9 34 139 12 54 251 67
15 47 11 35 149 21 55 257 69
16 53 10 36 151 19 56 263 72
17 59 12 37 157 15 57 269 81
18 61 8 38 163 16 58 271 79
19 67 6 39 167 24 59 277 85
20 71 11 40 173 18 60 281 83
Although the population increases considerably after stage 33, the data doesn’t provide too much confidence that it will continue to do so. We have to take a closer look at why this sequence is so erratic.
4. Analyzing the game
The size of the primes at each stage $s$ with the corresponding $p$, $p_{s}$ being the $s^{th}$ prime, is always about $1.25\cdot p_{s}\#$. The probability of a random number of this size being prime is $1/(\log p_{s}\#+0.23)$, or $1/(\theta(p_{s})+0.23)$ where $\theta(p)$ is the first Chebyshev function. For one prime $q$ of the stage s1s-1 there has to be on average at least one prime in the interval $q\cdot p_{s},(q+1)\cdot p_{s}]$ to have a chance that the sequence keeps on producing ever more primes. More precisely, the probability of getting at least one prime out of the respective interval must be bigger than getting no prime at all.
$\frac{\binom{p-1}{1}}{(\log q-1)^{1}}(1-\frac{1}{\log q})^{p-1}>\frac{\binom{p-1}{0}}{(\log q-1)^{0}}(1-\frac{1}{\log q})^{p-1}\>\text{ or simply }\>\frac{p-1}{\theta(p)-0.77}>1.$ (1)
This is true for $p\geq 3$ if and only if $\theta(p)<p-0.23$, or $\theta(p)<p$ for short, since the constant value $0.23=\log y$ is negligible for large $p$.
Now $\theta(p)$ is in fact smaller than $p$ most of the time — but only slightly so. And indeed $p-\theta(p)$ behaves just like the mercurial $Li(p)-\pi(p)$, along with the infinitude of sign changes (see Littlewood [7]):
Using the simple sum 111 As opposed to the “European convention” $Li(x)=\int_{2}^{p}\frac{\text{d}x}{\log x}$ and the “American convention” $Li(x)=\int_{0}^{p}\frac{\text{d}x}{\log x}$ which each differ from said sum by less than 1 for $p\geq 12.0050107$... (European) or $p\geq 2$ (American).
$Li^{*}(p)=\sum_{x=2}^{p}\frac{1}{\log x}\>,$ (2)
then
$p-\theta(p)=(\log p)(Li^{*}(p)-\pi(p))+1-\sum_{x=2}^{p-1}\bigg(\log(\frac{x+1}{x})(Li^{*}(x)-\pi(x))\bigg).$ (3)
This formula makes it also clear that $p-\theta(p)$ changes sign quite a while before $Li(p)-\pi(p)$ acquires negative values. For the sake of a ballpark figure: assuming $Li(x)-\pi(x)$ is on average close enough to $\frac{\sqrt{x}}{\log x}$, then $p-\theta(p)$ drops below zero by the time $Li(p)-\pi(p)<\frac{2\sqrt{p}+O(p^{1/3})}{(\log p)(\log p-2)}$. It can be expected that this happens for the first time close to the first $Li(x)-\pi(x)$ crossover near $1.4\cdot 10^{316}$ (Bays and Hudson [1], more extensive calculations by Saouter and Demichel [16]). The results of Platt and Trudgian [11] also confirm this.
Empirical data and heuristical reasoning suggests that $p-\theta(p)$ can usually be found in the vicinity of $\sqrt{p}$. The bias is given by Riemann’s prime counting formula [14]
$\pi(x)=\sum_{n=1}^{\infty}\frac{\mu(n)}{n}\bigg(Li(x^{1/n})-\sum_{\rho}Li(x^{\rho/n})+\int_{x^{1/n}}^{\infty}\frac{\text{d}u}{u(u^{2}-1)\log u}-\log(2)\bigg).$ (4)
If the Riemann Hypothesis (RH) is true, $Li(p)-\pi(p)$ oscillates along $\frac{\sqrt{p}}{\log p}$ with a magnitude of at most $O(\sqrt{p}\log p)$ (von Koch [17]), and so $p-\theta(p)$ oscillates along $\sqrt{p}$ with a magnitude of at most $O(\sqrt{p}\log^{2}p)$. Ingham [4] also proved this bias for sufficiently large $p$.
On a large scale, the number of intervals should increase slowly and asymptotically with $p-\theta(p)$, thus it may also decrease whenever $\theta(p)>p$. Yet the actual development strongly depends on the local irregularities of the distribution of prime numbers, so one can only assign the probability that the initial conditions are adequate.
Admittedly, (1) wasn’t quite accurate. The number of primes in an interval immediately after a semiprime as it is the case, especially in this short type of interval, is on average a bit smaller than for a randomly chosen interval. The following table illustrates this sort of difference for a small interval of ten numbers:
Table 4. For a random interval start and interval length 10, the average of numbers that are coprime to 2310=11#2310=11\# is above two, while for an interval of the same length after a prime number (or any number coprime to 11#11\#), significantly fewer than two such numbers are to be expected. Hardy and Littlewood [3] laid a solid foundation regarding this issue.
divisible by not interval
number 2 3 5 7 11 divisible sum
[…]
prime +1 100% 50% 25% 17% 10% 0% 0
prime +2 0% 50% 25% 17% 10% 28% 0.28
prime +3 100% 0% 25% 17% 10% 0% 0.28
prime +4 0% 50% 25% 17% 10% 28% 0.56
prime +5 100% 50% 0% 17% 10% 0% 0.56
prime +6 0% 0% 25% 17% 10% 56% 1.13
prime +7 100% 50% 25% 0% 10% 0% 1.13
prime +8 0% 50% 25% 17% 10% 28% 1.41
prime +9 100% 0% 25% 17% 10% 0% 1.41
[…]
Ultimately, the number of numbers that are relatively prime to k#k\# in the intervals in question $q\cdot p+1,(q+1)\cdot p-1]$ — denoted by $\Phi(p-1,k)$ — is
[…]
Yet disregarding that $p$ itself doesn’t appear as a factor in said interval, this particular formula is only valid for $2<k<p$. In contrast, for a random interval $R$, $\Phi_{R}(p-1,k)=(p-1)\cdot W(k)+O(1)$, where
$W(k)=\prod\limits_{\begin{subarray}{c}u=2\\ u\,prime\end{subarray}}^{k}\bigg(1-\frac{1}{u}\bigg).$ (6)
If $\Phi(p-1,k)$ is then divided by $W(k)\cdot\frac{p-2}{p-1}$ (as $p$ itself doesn’t appear as a factor in the intervals we’re looking at), then the result is an “adjusted” interval length. For this, $k$ must be appropriately large to attain the desired value ($k>\log p$, say).
There is a special connection between said adjusted interval length and the twin prime constant $C_{2}=0.6601618$…: for $k\to\infty$, the adjusted interval length $\Phi(p-1,k)/W(k)$, on the basis of a prime number preceding the interval, can be expressed as $C_{2}\cdot(p-1+a)$, with $a\in\mathbb{Q}$ being defined by
$a=2\cdot\sum_{x=1}^{\frac{p-1}{2}}\bigg(\prod\limits_{\begin{subarray}{c}r:\\ \text{every distinct}\\ \text{odd prime}\\ \text{factor of }x\end{subarray}}\frac{r-1}{r-2}-1\bigg).$ (7)
Some values of $a$ include
Table 5. The value of $\frac{a}{p-1}$ is asymptotic to $\frac{1}{C_{2}}-1-O(\frac{\log p}{2p\cdot C_{2}})$.
p1p-1 $a$ p1p-1 $a$
2 0 20 116/15
4 0 30 6866/495
6 2 40 141274/8415
8 2 50 1329632/58905
10 8/3 100 129132288244/2731483755
12 14/3 150 123443421975532168/1666745013838905
As $k\to\infty$, the adjusted interval length (for our case with the semiprimes) is — apart from $p=3,7,\text{ and }13$ — always a bit smaller than p1p-1. Dividing it again by $\log(q\cdot p)$ and we arrive at the expected average number of primes in one interval. The ratio of the adjusted interval length vs. p1p-1, which is equivalent to $\frac{\Phi(p-1,k)}{W(k)\cdot(p-2)}$ for appropriately large $k$, will be denoted below by $\psi(p)$.
Here we were focusing on heuristical aspects. For further reading about the structural analysis of the evolutional nature of the sequence, and potential vistas, we refer the reader to Santana [15] who gives an introduction on the notion of evolution on sets.
5. Proposing practical predictions
Using these heuristics, which are so far in very good agreement with the factual data, we can start to calculate the probabilities for the game to continue.
For example, there are 594 primes at stage 100, where $p=541$ and every prime $q=\lfloor p\#\cdot Y\rfloor$ has 220 decimal digits. Each $q$ then has a certain chance to yield $n$ primes in the following stages:
Table 6. Only the fittest will survive.
$s$ $p$ $n=0$ $n=1$ $n=2$ $n=3$ $n=4$ $n=5$ $n=6$ $n=7$ $n>7$
101 547 34.52% 36.75% 19.53% 6.90% 1.83% 0.39% 0.07% 0.01% 0.00%
102 557 49.78% 19.46% 14.17% 8.22% 4.33% 2.16% 1.03% 0.47% 0.36%
103 563 58.57% 12.14% 10.10% 7.05% 4.63% 2.95% 1.83% 1.12% 1.62%
104 569 64.34% 8.33% 7.46% 5.75% 4.22% 3.03% 2.14% 1.50% 3.22%
105 571 68.45% 6.11% 5.72% 4.68% 3.68% 2.84% 2.17% 1.64% 4.71%
110 601 78.69% 2.06% 2.14% 1.99% 1.80% 1.62% 1.44% 1.27% 8.99%
150 863 90.06% 0.10% 0.11% 0.11% 0.11% 0.11% 0.11% 0.11% 9.20%
200 1223 91.20% 0.01% 0.02% 0.02% 0.02% 0.02% 0.02% 0.02% 8.69%
1000 7919 91.48% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 8.52%
2000 17389 91.48% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 0.00% 8.52%
The column for $n=0$ is of importance to find out about the overall probability. (Once an interval is devoid of primes, it never comes back!) Within the first few thousand values of $s$, this value approaches $\omega\approx 91.476\%$. This gives the overall probability of $\omega^{594}\approx 1.039\cdot 10^{-23}$ that the sequence fails at this point.
Generally the value $\omega(s)$ can be defined for every $s\geq 1$ as the probability that a prime $q$ at stage $s$ will not cause an infinite sequence of primes in the further process.
For convenience, a recursive formula can be considered where, with $\psi(p)=\frac{\Phi(p-1,k)}{W(k)\cdot(p-2)}$ from above,
$\omega(s-1)=\bigg(1-\frac{\psi(p)}{\log q}\bigg)^{p-1}\sum_{j=0}^{p-1}\frac{\binom{p-1}{j}\omega(s)^{j}}{(\frac{\log q}{\psi(p)}-1)^{j}}$ (8)
and a starting point $\omega(t)=1-1/\sqrt{2p_{t}}$ for some $t$ can be set to quickly compute the values for all $s<t$. Varying $\omega(t)$ between 0 and $1-\varepsilon$ for some $\varepsilon>0$ leads to a lower and upper bound on $\omega(s)$. It may be useful that if $\omega(t)=\omega(s)+\delta$ then $\omega(t+1)\sim\omega(s+1)+\delta\cdot(1-\omega(s+1))/2$.
Again, we should bear in mind that these primes are not as flexible as (8) takes them to be: the probability of obtaining $n$ primes in the following step is zero when $n$ exceeds the maximal possible number of primes in the given interval (e.g. $n>97$ for p1=546p-1=546) as fixed in the specific rules for the k-tuple conjecture. To take account of that — at least to some degree —, $W(k)$ from (6) is deployed again such that k#<pk\#<p. For $p=547$, we set $k=7$ and $W(k)$ = 48/210 = 1/4.375. As $p$ gets larger than 11# = 2310, $W(k)$ can be changed to 480/2310 = 1/4.8125 and so on.
$\omega(s-1,k)=\bigg(1-\frac{\psi(p)}{W(k)\log q}\bigg)^{W(k)\cdot(p-1)}\sum_{j=0}^{W(k)\cdot(p-1)}\frac{\binom{W(k)\cdot(p-1)}{j}\omega(s,k)^{j}}{(\frac{W(k)\log q}{\psi(p)}-1)^{j}}$ (9)
This refined calculation reveals a slightly smaller value for $\omega(100,7)$ (compared to $\omega(100)$), namely 91.435%. The entire probability that the sequence fails at this point drops by 24% to $7.937\cdot 10^{-24}$. To show how vague these percentages are (especially at this rather early stage of computation), the actual portion of primes that don’t survive after stage 100 — at least 547 out of 594 — is $\geq 92.088\%$; to the $594^{th}$ power, this is over 68 times more than the value from the “refined” calculation! In other terms, 47 surviving primes as opposed to a predicted number of 51, from this point of view the deviation is not that bad (and it gets better for larger $p$).
$\omega(100)=\omega(100,1)=91.4760031\%$
$\omega(100,2)=91.4637729\%$
$\omega(100,3)=91.4515074\%$
$\omega(100,5)=91.4422850\%$
$\omega(100,7)=91.4345844\%$
It is difficult to pin down exact values there, but the refined calculation should at least give some indication on how large the error might be.
Not too surprisingly, $\omega(s)$ is largely situated in the neighborhood of $1-1/\sqrt{2p_{s}}$. This can be verified heuristically as follows: Proceed as to be seen in table 6, starting with stage $s$. Consider regular intervals with evenly distributed primes $q$, the size of the intervals being $z\cdot\log(q)$ for a chosen $z\geq 1$, so for lim $q\to\infty$, the probability that one interval doesn’t contain a prime at stage $s+1$ is $e^{-z}$. The probability to get 0 primes after $t$ steps at stage $s+t$ is $f(s+t)$, where $f(s)=0$ and recursively $f(x+1)=e^{z\cdot(f(x)-1)}$. As $\lim z\to 1$ for $s$ fixed, $x=1-2(z-1)+O((z-1)^{2})$, which can be approximated by $1-1/\sqrt{2p_{s}}$ for $z=p/\theta(p)\approx 1+1/\sqrt{p}$ (on average).
What is more, for $\theta(p)=p+m\cdot\sqrt{p}$, where $m=-1+O(log^{2}p)$ (under the RH),
$\omega(s+1)-\omega(s)\sim(\omega(s)-1)\bigg(\frac{m}{\sqrt{p}}+\frac{1}{p}+\frac{1-\omega(s)}{2}+O\bigg(\frac{\log p}{\sqrt{p^{3}}}\bigg)+O\bigg(\frac{1-\omega(s)}{\sqrt{p}}\bigg)\bigg)$ (10)
with error terms depending on either $p$ or $\omega(s)$. As long as $m\approx-1$ as is expected on average in the long run, if $\omega(s)$ is chosen a bit larger than $1-1/\sqrt{2p_{s}}$, the values for the subsequent $\omega(s+x)$ by the recursion formula above quickly head toward 1, so the first term in the parenthesis above, $m/\sqrt{p}$, becomes the most significant. If then at some point $Li(p)\approx\pi(p)$, thus $m\approx 0$, while $\omega(s)$ is very close to one, then the second term 1/$p$ becomes the most significant, but doesn’t nearly have the (then opposite) impact on $\omega(s)$ as in the region where $m\approx-1$. Now what does that say? If $\omega(s)$ stays close to 1, the initially chosen value is larger than the actual value, which is good because the actual value — the probability that the sequence terminates — is desired to be as small as possible. An $m$-value of +1 would adequately counter the effect of the expected average value 1-1, then again that should happen just as often as $m\leq-3$. It should be noted that there is yet no known effective region where $|m+1|\geq 2$, making it a candidate for a future endeavor. 222 Myerscough [9] tabulates the logarithmic density for the case $m>1$ as $1.603\cdot 10^{-95}$.
Once the number of primes $n$ — the number of possibilities for $y$ — reaches a “stable” value, one may give a rough estimate as to how the sequence continues. For instance, with $p$ = 1153 we have $n$ = 19690. In the next stage, with $p$ = 1163, one is likely to find approximately $n=19690\cdot\psi(p)\cdot(p-1)/(\theta(p)+0.23)\approx 20298$ primes.
Most of the time, the actual value lies between $n-\sqrt{n}$ and $n+\sqrt{n}$, and, by standard deviation measures, more than 99.7% of the time it should be well within $n\pm 3\sqrt{n}$. But even considering a negative deviation of $4\sqrt{n}$: $4\cdot\sqrt{20298}=-569.88$... is still outnumbered by the predicted growth 2029819690=60820298-19690=608, corroborating that the initial assumption holds with a very high probability if the distribution is normal (a sufficient but not particularly necessary condition).
Figure 1. Comparison. With the prime number output becoming more generous, the forecast is increasingly accurate.
Expanding this supposition, we can say the sequence is “stable” when $n>p$, where the average absolute growth rate $n/\sqrt{p}$ is bigger than the majority of the local deviations of order $\sqrt{n}$. This critical boundary is firmly overstepped at $s=98$: $p=521$ and $n=559$ (compared to $s=97$ with $p=509$ and $n=490$).
Looking at predicted vs. actual increase in terms of standard deviations, so far the sequence behaves pretty “normal”:
[…]
Long-term predictions imply that the number of primes may surpass $10^{6}$ at $\approx$ stage 339 ($p=2281$), $10^{9}$ at $\approx$ stage 702 ($p=5297$), $10^{12}$ at $\approx$ stage 1210 ($p=9811$), $10^{100}$ at $\approx$ stage 118948 ($p=1568341$), and may have reached $n\approx 10^{10^{155}}$ by the time that $\theta(p)>p$ where we expect to see the next slight decline of the sequence ever since stage 139 (see table 7).
The following seemingly unrelated approximation formula for $n$ can be used in the medium term:
$n\approx\frac{\sqrt{e^{\sqrt{3s}}}}{8}$ (11)
which is surprisingly accurate in spite of its simplicity in the narrow sense that it is off by less than 20, with one little exception, for $s\leq 61$, off by at most 20% for $98\leq s\leq 114$, and off by at most 15.4% for $115\leq s\leq 331$ — compared to more elaborate predictions, the latter may hold even for a couple more $s\leq 500$. Furthermore, (11) may be off by a factor of no more than 2 for $66\leq s\leq 1000+$. In the long run, however, the expression in (11) is expected to far exceed $n$, thus calling into question any more daring predictions based on it.
6. The status quo, pt. I: risk of failure
At stage 331 ($p=2221$) there are 810387 probable primes, and $\omega(331)\approx 95.45431\%$. $0.9545431^{810387}\approx 3.71\cdot 10^{-16374}$ — that is, for the time being, the conjectured probability that the game fails for some $s>331$.
Up to that point, 32103994 numbers have not yet been tested for certified primality, the smallest of those being in the ballpark of $2.38\cdot 10^{247}$. No counterexample is known for the combination of a Baillie-Pomerance-Selfridge-Wagstaff pseudoprimality test and a strong Lucas test which were carried out on the probable primes. 333 including some trial division, additionally decreasing the chances that a pseudoprime slipped through However if there is one among all the established PRPs, it could take effect on at most 85448 PRPs (the most prolific branch at $p=601$). $0.9545431^{810387-85448}\approx 1.00\cdot 10^{-14647}$, i.e. $2.70\cdot 10^{1726}$ times larger. It is yet unclear how much this will affect the above evaluation, especially by taking a look at Pomerance [13] who gives a heuristic argument that the number of counterexamples for a BPSW primality test up to $x$ is $\gg x^{1-\varepsilon}$ for some $\varepsilon>0$ and sufficiently large $x$. In either case, said ratio can be systematically reduced by checking the primality of the prolific branches in the sequence.
7. The status quo, pt. II: closing in on $y$
Now, what about the initially proposed values for $y$? The lower and upper bounds are easily obtained by picking the smallest prime $q_{s,1}$ and the largest prime $q_{s,n}$ of the last computed stage and dividing by p#p\#: $y_{\min}=q_{s,1}/p\#$ and $y_{\max}=(q_{s,n}+1)/p\#$. The calculations so far demonstrate that any $y$ lies in a very small range between
1.2541961015780119362776795549142134237798692180426221958327225546088646994287514475
and
1.2541961015780119362776795549142134237798692180426221958327225546088646994290445894.
In particular, $y_{\max}-y_{\min}=2.931419...\cdot 10^{-76}$.
Although it is not dead certain that such a $y$ exists, if it actually does, then there should be infinitely many of those. Otherwise, we’d have to assume a unique solution where the probability of getting one prime for $p\to\infty$ — the column for $n=1$ in Table 6 — is greater than zero. But if this is the case, then at every stage a factor of approximately $1/e$ falls back on the column for $n=0$ which then would converge to 1, challenging our assumption that there is a unique solution. It would be nice, however, to have a rigorous proof.
8. Ancestors and survivors
Having calculated all primes up to a certain stage, we can trace back these primes for previous stages and see which of those have “survived” in the long run. All probable primes from stage 331 originated from one prime at stage 43. Recall that there were a total of 14 primes at stage 43, so the other 13 “petered out” along the way, including one branch that originated in stage 37 tenaciously keeping up until it succumbed at stage 95.
The first split then appears to be at stage 44. From there on, at least two values of $y$ may satisfy the hitherto harsh conditions. And suddenly there’s a lot more to come:
Figure 3. Number of possible values for $y$ (for all we know). From $s=95$ onward, all primes $q$ emerge from one common ancestor at $s=43$, namely 1292942159746921794791923187781692727375711831825607985864285936838920812561. Naturally, the branches in the picture are not equally strong. For $s=331$ ($p=2221$), the following number of probable primes divide up on the nine given groups: 11477 + 9305 + 64706 + 123386 + 137475 + 101078 + 13648 + 77384 + 271928.
If one was to specify these values as $y_{0}$, $y_{1}$ and so on to attain fixed constants for the various possibilities, one might set $y_{0}$ as the minimum value $y_{min}$ as described above, and label each of the consecutive numbers $y_{1}$, $y_{2}$ … as the other smallest possible value(s) after a precalculated split, when it is “safe enough to assume” that those branches are stable.
So $y_{0}=y_{min}$ (see above), and then
$y_{1}=y_{0}+2.9314187260027917698243…\cdot 10^{-76}$
(from the first split at stage 44, $\lfloor y_{0}\cdot 191\#\rfloor\cdot 193+30$ and +88+88, $y_{1}\approx y_{0}+58/193\#$)
$y_{2}=y_{0}+1.9278096567063412938136…\cdot 10^{-79}$
$y_{3}=y_{2}+1.1844960407807006213569…\cdot 10^{-80}$
$y_{4}=y_{2}+1.5406827193677358123185…\cdot 10^{-80}$
$y_{5}=y_{0}+1.5619786064199049976025…\cdot 10^{-100}$
$y_{6}=y_{0}+1.4764435119323593401152…\cdot 10^{-104}$
$y_{7}=y_{6}+1.4210514603041973809966…\cdot 10^{-111}$
$y_{8}=y_{1}+2.5503871104036223556538…\cdot 10^{-119}$
The value $y_{6}$ is one that may be doubted more than $y_{0}$$y_{5}$ since it derives from the weakest branch of Figure 3 with only 9305 primes. But since 9305 >> 2221 (thus a stable branch), it can be expected to remain valid.
The stability criterion as described above, $\lim_{sup}n/p<1$ for finite branches, may even be strengthened to $\lim_{sup}n/({\frac{1}{2}}\cdot\sqrt{p}\cdot\log p)<1+\varepsilon$ and still hold for the entirety of the sequence for relatively small $\varepsilon$. Even if this is not true for all $\varepsilon>0$ (and under certain conditions — looking at runs of consecutive non-surviving primes of a given stage — can be anticipated to fail), in practice, it will be a safe bet that no-one will explicitly find a counterexample for branches with more than $\frac{1}{2}\cdot\sqrt{p}\cdot\log p$ primes. There might be a bound $\lim_{sup}n/({\frac{1}{2}}\cdot\sqrt{p}\cdot\log p)\leq c$ for finite branches for some constant $c$, and thus far $c>0.75478$ ($177^{th}$ branch/prime of stage 111 with 64 primes at stage 123 but dissipating at stage 316). So two more questions are probably to remain unanswered: Is there such a bound, maybe with $c>1$? If yes, what is the heuristically/actually largest $c$? If no, is there any other bound?
The bifurcation tree as depicted in Figure 3 gets quite impressive when expanded. For example, tracing back the primes from stage 331 to stage 100 crystallizes out 47 primes (compared to originally 594) with an eventful history.
Figure 4. Alleged number of possibilities up to stage 100.
It is interesting to note the threefold split at stage 46, $p=199$. If all of these primes persist an infinite run, this still is the only point for $s\leq 100$ where more than two primes out of a single prime from the previous stage will simultaneously continue the race. But how probable is this now? We have to consider the possible number of primes arising from one prime of the previous stage and the possibility that at least three of them become stable sequences, including the permutations if more than three primes show up simultaneously in the current stage:
$\sum_{x=3}^{\infty}\frac{\binom{p-1}{x}}{(\log q-1)^{x}}(1-\frac{1}{\log q})^{p-1}(1-\omega(s-1))^{x}\binom{x}{3}$ (12)
For $s=101$, this value is 0.0043%, thus $\approx$ 2.76% for all 594 primes of stage 100.
Figure 5. The value is growing in the process (e.g. about 52% for $s=252$), so we can soon expect another threefold split. In fact, the next one appears to be in the upper region of $p=613$, $s=112$.
Substituting 3 for $k$ in (12), it can handily be used for the probability of a $k$-fold split for any $k\geq 2$.
The possibly last stage without any split, that is without a temporarily increasing number of surviving primes and hence admissible values for $y$, is found at $s=103$ ($p=563$).
9. Broader utilization
It’s not without a little luck that the sequence as described exists. If we look for $y$’ such that $y’\cdot p\#$ rounded to the nearest integer is prime, we wouldn’t get far: at stage 23, one prime would be left: $\lfloor y’\cdot 83\#+\frac{1}{2}\rfloor=q’_{23,1}=206780313999369083332356327764879$. There’s no prime within $q’_{23,1}\cdot 89\pm 44$, so that sequence would be finite (or “mortal”) — even though the primes are smaller compared to $y=1.2541961$... ($q’_{s}=\{2,2,2,2,1,1,2,3,6,4,1,3,5,4,2,2,$ 2,4,5,5,2,2,1,0.}2,4,5,5,2,2,1,0.\} — and the ratio $\frac{y}{y’}=1.6198$... is close to the golden ratio:)
And for those who can’t get enough sequences that grow like $y\cdot p\#$, here is a definition for “semi-p#Y-sequences”: find $y$ such that $\lfloor y\cdot p\#\rfloor$ is a prime number for every $p\geq 3$. The only difference here is that $\lfloor 2y\rfloor$ is allowed to be an even number other than 2 (maybe even a semiprime, depending on the taste about restrictions, though it’s hard to find such an example like $\lfloor y\rfloor=210457743323$ — with 1124 primes at $s=331$). This leaves a lot of sequences on the verge of extinction for quite a while during the calculations. Some of them should survive an infinite run:
$0<y<1$: this gives a sequence which is stronger than the original p#Y-sequence
($1<y<2$: original sequence)
$25<y<26$: stronger than the original sequence up to stage 166, weaker after stage 168
$1411<y<1412$: a weak sequence with only 883 primes at $s=168$, $p=997$
$3432<y<3433$: a very weak sequence, only 168 primes at $s=168$, $p=997$
$13948<y<13949$: another weak sequence
$201420<y<201421$: another one which has somewhat good chances to survive
$6007103<y<6007104$: rather weak
$25510020<y<25510021$: just a bit stronger than $\lfloor y\rfloor=3432$ until $s=167$, $p=991$
… Can you find others? Not like these, they’re only finite (successive records):
$2<y<3$: extinct at stage 4
$3<y<4$: extinct at stage 7
$5<y<6$: extinct at stage 11
$16<y<17$: extinct at stage 14
$978<y<979$: extinct at stage 17
$6640<y<6641$: extinct at stage 23
$11456<y<11457$: extinct at stage 35
$160563<y<160564$: extinct at stage 38
$283257<y<283258$: extinct at stage 68
$1117230<y<1117231$: extinct at stage 78
$1594501<y<1594502$: extinct at stage 86
$55990660<y<55990661$: extinct at stage 106
$108286142<y<108286143$: extinct at stage 114
The number of $y$’s with different integer part that result in infinite semi-sequences should also be infinite.
10. Appendix A. More on the evolution
Table 7. Data for $s\leq 330$, using BPSW-pseudoprimality for the primes $q$ for $s\geq 110$. nn* stands for the number of surviving branches, backtracked from the 810387 primes of stage 331. This number seems to be settled for $s\leq 130$ but will be subject to change for $s\geq 131$ by the time the calculation is taken further since many of them will get cancelled out, so this column only gives a momentary picture (as well as an upper bound). Nevertheless, the rate of decrease in nn* for larger $s$ may be of interest in itself.
$s$ $p$ $n$ $s$ $p$ $n$ nn* $s$ $p$ $n$ nn* $s$ $p$ $n$ nn*
1 2 1 31 127 2 1 61 283 93 8 91 467 413 35
2 3 1 32 131 5 1 62 293 81 9 92 479 426 36
3 5 1 33 137 6 1 63 307 81 9 93 487 446 37
4 7 1 34 139 12 1 64 311 67 9 94 491 454 37
5 11 2 35 149 21 1 65 313 66 11 95 499 491 37
6 13 2 36 151 19 1 66 317 74 11 96 503 456 38
7 17 3 37 157 15 1 67 331 91 12 97 509 490 38
8 19 3 38 163 16 1 68 337 88 12 98 521 559 41
9 23 4 39 167 24 1 69 347 90 14 99 523 573 43
10 29 6 40 173 18 1 70 349 95 15 100 541 594 47
11 31 4 41 179 18 1 71 353 102 16 101 547 652 47
12 37 5 42 181 17 1 72 359 126 18 102 557 718 50
13 41 5 43 191 14 1 73 367 152 19 103 563 757 50
14 43 9 44 193 24 2 74 373 154 19 104 569 786 53
15 47 11 45 197 24 3 75 379 166 19 105 571 835 55
16 53 10 46 199 28 5 76 383 187 20 106 577 839 57
17 59 12 47 211 30 5 77 389 214 20 107 587 854 59
18 61 8 48 223 36 5 78 397 206 21 108 593 906 63
19 67 6 49 227 49 5 79 401 201 21 109 599 916 65
20 71 11 50 229 44 5 80 409 220 21 110 601 998 69
21 73 5 51 233 52 5 81 419 241 23 111 607 988 71
22 79 4 52 239 53 5 82 421 249 25 112 613 1016 74
23 83 6 53 241 55 5 83 431 269 25 113 617 1078 78
24 89 3 54 251 67 6 84 433 320 27 114 619 1165 81
25 97 2 55 257 69 6 85 439 354 29 115 631 1237 84
26 101 1 56 263 72 7 86 443 354 31 116 641 1295 86
27 103 3 57 269 81 7 87 449 365 32 117 643 1371 90
28 107 1 58 271 79 7 88 457 369 33 118 647 1399 93
29 109 1 59 277 85 8 89 461 358 33 119 653 1523 97
30 113 3 60 281 83 8 90 463 387 34 120 659 1618 101
Stage 139 is the last known point where the number of primes decreases. (Interestingly enough, just before that happens, we have $n=4p$, which is the only known — but certainly not the only — example where $n$ is a multiple of $p$.)
$s$ $p$ $n$ nn* $s$ $p$ $n$ nn* $s$ $p$ $n$ nn*
121 661 1701 105 156 911 5971 387 191 1153 19690 1138
122 673 1773 108 157 919 6366 401 192 1163 20364 1169
123 677 1801 116 158 929 6582 411 193 1171 21154 1205
124 683 1889 121 159 937 6799 430 194 1181 21954 1251
125 691 1885 127 160 941 7006 452 195 1187 22688 1274
126 701 1944 132 161 947 7151 460 196 1193 23528 1319
127 709 2008 138 162 953 7599 471 197 1201 24295 1349
128 719 2141 144 163 967 7920 491 198 1213 25085 1387
129 727 2136 150 164 971 8276 508 199 1217 25717 1428
130 733 2241 158 165 977 8708 521 200 1223 26680 1477
131 739 2354 165 166 983 9086 532 201 1229 27569 1510
132 743 2442 177 167 991 9577 550 202 1231 28242 1555
133 751 2548 179 168 997 9736 566 203 1237 28917 1595
134 757 2649 184 169 1009 10329 581 204 1249 29871 1639
135 761 2791 193 170 1013 10722 589 205 1259 30707 1688
136 769 2998 200 171 1019 11048 606 206 1277 31555 1733
137 773 3017 207 172 1021 11387 634 207 1279 32633 1786
138 787 3148 214 173 1031 11769 663 208 1283 33494 1849
139 797 3132 224 174 1033 12251 679 209 1289 34556 1899
140 809 3385 231 175 1039 12490 705 210 1291 35744 1957
141 811 3537 245 176 1049 12905 727 211 1297 36546 2010
142 821 3698 250 177 1051 13209 743 212 1301 37363 2063
143 823 3914 261 178 1061 13556 768 213 1303 38018 2123
144 827 4175 266 179 1063 13918 790 214 1307 38936 2188
145 829 4345 273 180 1069 14393 815 215 1319 39633 2272
146 839 4360 284 181 1087 15006 845 216 1321 40152 2344
147 853 4537 291 182 1091 15472 867 217 1327 40822 2420
148 857 4722 303 183 1093 15695 901 218 1361 42524 2500
149 859 4862 307 184 1097 16075 930 219 1367 44138 2580
150 863 4951 319 185 1103 16548 963 220 1373 45674 2654
151 877 5099 331 186 1109 16852 999 221 1381 47308 2738
152 881 5228 344 187 1117 17548 1024 222 1399 49203 2818
153 883 5334 356 188 1123 17926 1055 223 1409 50999 2905
154 887 5460 363 189 1129 18415 1081 224 1423 53578 2982
155 907 5715 377 190 1151 19145 1115 225 1427 56009 3071
$s$ $p$ $n$ nn* $s$ $p$ $n$ nn* $s$ $p$ $n$ nn*
226 1429 58245 3143 261 1663 141747 8077 296 1949 323493 26835
227 1433 60614 3264 262 1667 144129 8290 297 1951 332369 28016
228 1439 63012 3359 263 1669 146278 8514 298 1973 344077 29288
229 1447 65622 3434 264 1693 150011 8752 299 1979 355840 30683
230 1451 67609 3542 265 1697 153719 9016 300 1987 367371 32050
231 1453 69411 3650 266 1699 157438 9268 301 1993 379303 33504
232 1459 71677 3745 267 1709 160935 9582 302 1997 391133 35062
233 1471 73978 3854 268 1721 164713 9877 303 1999 402306 36665
234 1481 76816 3964 269 1723 168209 10187 304 2003 413306 38433
235 1483 78753 4049 270 1733 171914 10497 305 2011 424348 40351
236 1487 81051 4156 271 1741 175208 10838 306 2017 435084 42411
237 1489 82647 4260 272 1747 178924 11177 307 2027 446383 44703
238 1493 84106 4386 273 1753 182442 11549 308 2029 457067 47035
239 1499 85756 4511 274 1759 186286 11951 309 2039 468311 49593
240 1511 87469 4631 275 1777 190466 12354 310 2053 482068 52470
241 1523 90333 4765 276 1783 196294 12773 311 2063 494789 55576
242 1531 93284 4879 277 1787 200525 13184 312 2069 509095 58953
243 1543 96279 5012 278 1789 204944 13625 313 2081 524247 62650
244 1549 99150 5141 279 1801 209420 14097 314 2083 538695 66722
245 1553 101723 5281 280 1811 214672 14574 315 2087 551951 71192
246 1559 104560 5402 281 1823 218990 15100 316 2089 564045 76342
247 1567 107041 5542 282 1831 224272 15623 317 2099 577252 81969
248 1571 109795 5691 283 1847 230486 16200 318 2111 592642 88171
249 1579 111912 5827 284 1861 238339 16792 319 2113 605293 95404
250 1583 114508 5989 285 1867 246127 17405 320 2129 620280 103655
251 1597 117842 6160 286 1871 253236 18037 321 2131 635284 113441
252 1601 121087 6327 287 1873 259946 18726 322 2137 649598 124938
253 1607 124407 6493 288 1877 266707 19481 323 2141 664112 138630
254 1609 126382 6682 289 1879 272114 20242 324 2143 676043 155455
255 1613 128949 6854 290 1889 278512 21025 325 2153 689579 176551
256 1619 130732 7021 291 1901 285821 21851 326 2161 703065 203256
257 1621 132088 7226 292 1907 292121 22714 327 2179 720414 240371
258 1627 133678 7440 293 1913 298404 23666 328 2203 742364 292120
259 1637 135308 7642 294 1931 305779 24693 329 2207 763990 370783
260 1657 137970 7862 295 1933 314060 25728 330 2213 787038 506979
Data for $s=331$ as output from the program on the following page is given as an ancillary file to this version of the paper.
[…]
a=350842542483891235293716663559065020274899073; d=[0];
s=0; b=a;
while(b>1, s++; b\=prime(s));
i=#d; gettime();
while(s<330,
   s++; p=prime(s); o=a*p; c=d;
   e=99+floor(i*(1+2/sqrt(p))); d=vector(e);
   m=i; i=0;
   for(j=1, m,
      print1("interval "j"/"m,Strchr(13));
      o+=c[j]*p; v=vector(p);
      forprime(b=3, p-2, r=b-lift(Mod(o,b)); forstep(l=r, p, b, v[l]=1));
      forprime(b=p+2, p^2, r=b-lift(Mod(o,b)); if(r<p, v[r]=1));
      forstep(r=2, p-1, 2,
         if(!v[r],
            q=o+r;
            if(ispseudoprime(q),
               i++; if(i>1, d[i]=q-z, a=q); z=q;
               print1("interval "j"/"m" - "i" primes found",Strchr(13))
   g=floor(gettime()/1000); x="[CPU time: ";
   f=floor(g/3600); if(f,x=Str(x,f"h "));
   f=floor(g/60); if(f,x=Str(x,f%60"m "));
   x=Str(x,g%60"s] "); t=Str("Stage "s); print(t" (p="p"): "i" primes "x);
   t=Str("p#Y "t".txt"); write(t,"a="a"; d="vecextract(d,Str("1.."i)))
[…]
11. Appendix B. “Mille”stone
The first titanic p#Y-prime ($s=350$, $p=2357$, primality proved over all stages): $q_{350,1}=$ 297588802944669004567432988935086641892218253950808346654606799666416703403235386 110626587476071000262691671679224513211575176049294913346932056675348621165796730 263895561614248056123980950034528686832211500014435248604833958209819278373161550 816949962259892736268616001731340442751214329425997049930689376917146879956304444 951698385088539246081531976376843432978119495609784805439847386853846175609773406 294780298954522227295543214788475948953215376258100904758911480647631719489690751 598932601226480640025811451192259365414314027908668601256915767355990441760663317 227840259709936806311906508957316674862617637555729680550552499945756091942766728 462111799759803946513079633400975724089109421146044040340589262506167584698451951 787089619451193713787721236181896575926339065362406372104874478640876833784201015 302057350287221183373441732847205339141787284034897239162684632744754226872565733 662978724742861772240754047058150049733225603812946433723077381838542363028349647 7698386126623480539308646019
It’s a nice fact that 2357# of all primorials has exactly 1000 digits.
The ostensibly first titanic p#Y-prime that survives in the long run, the $16^{th}$ smallest of them all at $p=2357$, spanning its branch beyond 13751# ($s>1626$), is $q_{350,1}+$ 648890029246175571592953681722255270285895737015173251749453742478642415683875459 622820131990229693513724200589337798081947713889805053628824480433575300758724935 31089513889637480426037406091083077809690,
which leads to $y_{min}\geq$
1.25419610157801193627767955491421342377986921804262219583272255460886469942875144 751323169673647331200713029313835829410519055068071464454595777347989721472473549 850870383774045538648544910432104569800625071405146132797606019221734399136669410 231685026270656941987419822020206973004280705089121259525801855611303610702614940 065145204852907234077802800854431673412825358322402127785595190344370611325045336 730133684806271064030484811296434802163878424799778915265485948593800601469382480 980040875271653561997759928697347239814789283725106267672399076512871960330242356 126907949160683090804262705523408008158909911209586208398778315843260104151172880 387141761635412295558476482556483533650156157283843179063498926572476414145174204 072485693292701787200666665412458613015446492323282769411572687097469909612014411 743405605429162630100500385105709425493894716245564452834943026368056855139809017 404605931009796839371601349304261945450753825556512381095336296655187077486338594 1160720014897114954349144158…
The first 129 primes of stage 428 ($p=2969$) are known to be “mortal”: $q_{428,130}$ is the first that may survive an infinite run, which is a currently known maximum. It takes on average about $\sqrt{p}/2$ primes to find a surviving prime. An error term for this approximation, assumed to be of order of at most $O(\sqrt{p}/\log p)$, would probably be desireable.
12. Appendix C. Sequence tuples (multiple primes in one interval)
First twin: ($1^{st}$ prime of stage 4)*11+ {4, 10}
First triplet: ($1^{st}$ prime of $s$ = 7)*19+ {10, 12, 16}
First quadruplet: ($2^{nd}$ prime of $s$ = 9)*29+ {6, 14, 24, 26}
First quintuplet: ($18^{th}$ prime of $s$ = 46)*211+ {42, 110, 140, 144, 194}
First sextuplet: ($69^{th}$ prime of $s$ = 69)*349+ {18, 40, 234, 262, 292, 298}
First septuplet: ($238^{th}$ prime of $s$ = 84)*439+ {34, 228, 252, 282, 364, 378, 382}
First octuplet: ($5687^{th}$ prime of $s$ = 159)*941+ {170, 234, 294, 462, 696, 740, 752, 812}
First nonuplet: ($270^{th}$ prime of $s$ = 181)*1091+ {18, 46, 60, 166, 180, 312, 850, 1062, 1080}
First decuplet: ($193057^{th}$ prime of $s$ = 289)*1889 + {220, 238, 378, 624, 934, 1048, 1414,
. 1612, 1678, 1750}
Higher-order descendants (first instances of number of “grandchildren”):
3 descendants of order 2: ($1^{st}$ prime of stage 6)*17*19+ {162, 164, 168}
4 descendants: ($1^{st}$ prime of $s$ = 7)*19*23+ {246, 284, 384, 386}
5 descendants: ($5^{th}$ prime of $s$ = 21)*79*83+ {3504, 3520, 5190, 5200, 5224}
6 descendants: ($4^{th}$ prime of $s$ = 12)*43*47+ {102, 108, 582, 598, 1428, 1450}
7 descendants: ($7^{th}$ prime of $s$ = 34)*149*151+ {754, …, 12484}
8 descendants: ($30^{th}$ prime of $s$ = 55)*263*269+ {7686, …, 40560}
9 descendants: ($10^{th}$ prime of $s$ = 59)*281*283+ {11388, …, 59016}
10 descendants: ($16^{th}$ prime of $s$ = 47)*223*227+ {13650, …, 38340}
11 descendants: ($42^{nd}$ prime of $s$ = 67)*337*347+ {24432, …, 116892}
12 descendants: ($27^{th}$ prime of $s$ = 55)*263*269+ {45742, …, 68590}
13 descendants: ($3053^{rd}$ prime of $s$ = 138)*797*809+ {195918, …, 603780}
14 descendants: ($623^{rd}$ prime of $s$ = 101)*557*563+ {94812, …, 261562}
15 descendants: ($32^{nd}$ prime of $s$ = 72)*367*373+ {2424, …, 134612}
16 descendants: ($45231^{st}$ prime of $s$ = 224)*1427*1429+ {74772, …, 1944498}
17 descendants: ($2030^{th}$ prime of $s$ = 176)*1051*1061+ {74496, …, 1083072}
18 descendants: ($376869^{th}$ prime of $s$ = 321)*2137*2141+ {1058496, …, 3727156}
The like-minded reader with enough time on their hands might be encouraged to find successive maxima of descendants of higher order. To start, here’s 4 descendants of order 3:
($1^{st}$ prime of stage 6)*17*19*23+ {3742, 3780, 3880, 3882}
13. Appendix D. Relative strength of split intervals
Figure 6. A nice and colorful edit showing the strength of the first three split intervals, starting at stage 45. The 592642 probable primes at stage 318 are here divided into three groups (154145 + 184653 + 253844) with the respective relative strength 26.01%, 31.16%, and 42.83%. These values may (hopefully) stabilize while the number of primes in the sequence continues to grow. Giving the exact percentages as $s$ goes to infinity will be rather difficult.
14. Appendix E. Plouffe, revisited
As a final tidbit of numerical info, inspired by Plouffe’s work mentioned in the introduction, we give a constant $A$ (to 1029 places after the decimal point) such that $\lfloor A^{1.001^{n}}\rfloor$ should give an infinite sequence of primes, for integer $n\geq 0$:
$A=10^{3875}$- 9840440.013269045402699925892590230034000537023543058287831784100957 276330984498140336774521413625639037370179638123065989644921831945828284474797146 214314469465943653321876212421217357462243438103894210660374222702124802145689793 662592212239695486359151401226648141442872210231891779023391189805301737658502412 189281932141114052020391866640443434821779837126175086985360497145764015143114697 436609454475477982979703079133332121473465799127802190473213691503591043455180286 502985352685402044844822486579962417722857237662143660192493843854713085865446265 026287773151577320009216914265110643337790882410603483306933710003501741787035605 297213202524866787149107203669351182374236682986886760407222192912257430323369703 769706007229005048714432375371543157174405792236627110947314049230460840453363435 080982900105891219946283712237437066606879315185506382375652998319912097105539735 706790752939943728420192710162069587039287823862485223550170317760285385396844814 481707064610195932400139047255566802623529101069361896409968571400144014982338…
We would be happy to learn about a smaller known value for $A$.
15. Acknowledgements
[…]